# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v1

@@ 6. Computing isomorphism classes admitting a polarisation of degree $4$
In [18], it is described how to compute the isomorphism classes of abelian varieties over $\mathbb{F}_{q}$ belonging to an ordinary isogeny class $\mathcal{C}$ determined by an irreducible Weil polynomial $f(t)$.
Let $K=\mathbb{Q}[t]/(f(t))=\mathbb{Q}[\pi]$ where $\pi$ denotes the class of $t$ in $K$ and set $R=\mathbb{Z}[\pi,q/\pi]$. We fix once and for all an isomorphism ¯$\overline{\mathbb{Q}}_{p}\simeq\mathbb{C}$, which induces a $p$-adic valuation $v_{p}$ on the complex numbers. Let $\Phi$ be a CM-type of $K$, that is, a choice of half of the homomorphisms $K\to\mathbb{C}$, one per conjugate pair. Assume furthermore that $\Phi$ satisfies the Shimura-Taniyama formula, that is, $v_{p}(\varphi(\pi))>0$ for every $\varphi\in\Phi$. We say that $\lambda\in K$ is $\Phi$-positive if $\Im(\varphi(\lambda))>0$ for every $\varphi\in\Phi$.
We summarise here the results that are relevant for us; see [18, Cor. 4.4, Thm. 5.4].
  • There is a functor $\mathcal{F}$ from the category of abelian varieties in $\mathcal{C}$ (with $\mathbb{F}_{q}$-morphisms) and the category of fractional $R$-ideals in $K$ (with $R$-linear morphisms).
  • If $\varphi:A\to B$ is an isogeny in $\mathcal{C}$, $\mathcal{F}(A)=I$ and $\mathcal{F}(B)=J$ then $\mathcal{F}(\varphi)$ is multiplication by an element $\lambda\in K^{\times}$ such that $\lambda I\subseteq J$.
  • If $\mathcal{F}(A)=I$ then $\mathcal{F}(A^{\vee})=\overline{I}^{t}=\{z\in K:\mathop{\rm Tr}\nolimits_{K/\mathbb{Q}}(zI)\subseteq\mathbb{Z}\}$.
  • Given an isogeny $\lambda:A\to A^{\vee}$, set $\mu=\mathcal{F}(\lambda)$. Then $\lambda$ is a polarisation if and only if $\lambda=-\overline{\lambda}$ and $\lambda$ is $\Phi$-positive. Moreover, $\deg(\varphi)=|\overline{I}^{t}/\lambda I|$.
  • If $\lambda$ and $\lambda^{\prime}$ are two polarisations of the same abelian variety $A$, then the polarised abelian varieties $(A,\lambda)$ and $(A,\lambda^{\prime})$ are isomorphic if and only if $\mathcal{F}(\lambda)=v\bar{v}\mathcal{F}(\lambda^{\prime})$ for some unit $v\in S^{\times}$ where $S=(\mathcal{F}(I):\mathcal{F}(I))$ is the endomorphism ring of $A$.
Given the previous considerations one obtains an algorithm to compute the abelian varieties in $\mathcal{C}$ together with their polarisations of a fixed degree, up to polarised isomorphism. See [18, Sec. 6] for more details. We will use this algorithm in the next examples.
@@
Consider the isogeny class of abelian surfaces over $\mathbb{F}_{2}$ with label 2.2.a_ab determined by the Weil polynomial $f(t)=t^{4}-t^{2}+4$. According to the LMFDB, the class contains a Jacobian. Moreover, one computes that the order $\mathbb{Z}[\pi,4/\pi]$ has $3$ overorders $S_{1}$, $S_{2}$ and $\mathcal{O}$, where $\mathcal{O}$ is the maximal order of the number field $K=\mathbb{Q}[t]/(f(t))$. One observe that $S_{2}=\overline{S_{1}}$. Hence, each abelian surface $A$ with endomorphism ring $S_{1}$ cannot be isomorphic to its dual, which has endomorphism ring $S_{2}$. In particular, such an $A$ does not admit a principally polarisation. Hence, such an $A$ does not contain a curve of arithmetic genus $2$, as explained in Remark 1.3, even if it is isogenous to a Jacobian.
@@
In Theorem 5.7, we see that, in the isogeny class $\mathcal{C}$ associated to an ordinary Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, there is an abelian variety $A$ admitting a polarisation of degree $4$ if and only if there is an abelian variety $A^{\prime}$ in $\mathcal{C}$ with maximal endomorphism ring admitting a polarisation of degree $4$. This is not the case in general. For example, in the isogeny class with label 2.13.a_al determined by $t^{4}-11t^{2}+13^{2}$ there are abelian varieties admitting a polarisation of degree $4$ by none of them has maximal endomorphism ring.
@@ 7. Genus $3$ curves lying on abelian surfaces
In this final section, we switch our attention from abelian surfaces to curves, and gather information on genus $3$ curves lying on abelian surface with Weil polynomials in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, that is, which is simple and not isogenous to a Jacobian, or equivalently, not containing curves of geometric genus 00,$1$ or $2$. We fix an abelian surface $A$ defined over $\mathbb{F}_{q}$ in such an isogeny class, and we suppose that there is an absolutely irreducible projective smooth genus $3$ curve $C$ defined over $\mathbb{F}_{q}$ that lies on $A$. The following lemma characterises the Jacobian of $C$.
@@
Let $q$ be the power of an odd prime. Let $C$ be a genus $3$ curve lying on an abelian surface $A$ defined over $\mathbb{F}_{q}$. Then $C$ is the double cover of an elliptic curve $E$ and the Jacobian $\mathop{\rm Jac}\nolimits(C)$ of $C$ is isogenous to $E\times A$.
@@
By [5, Prop. 1.8], as $C$ lies on $A$, it is bielliptic, that is, it is the double cover of an elliptic curve $E$. In particular, we know that the Jacobian of $C$ is isogenous to $\mathop{\rm Jac}\nolimits(E)\times P\simeq E\times P$, $P$ being the Prym variety associated with the double cover. Hence, since by [5, Sec. 1.4] we know that $P$ and $A$ are isogenous, we deduce $\mathop{\rm Jac}\nolimits(C)\sim E\times A$. ∎
We recall that absolutely irreducible genus $3$ curves are either hyperelliptic curves or plane quartics. The next lemma shows that hyperelliptic genus $3$ curves cannot lie on our abelian surfaces, at least if we work over finite fields of odd characteristic.
@@
Let $q$ be the power of an odd prime. Let $C$ be a genus $3$ curve lying on an abelian surface $A$ defined over $\mathbb{F}_{q}$ which is simple and not isogenous to a Jacobian. Then $C$ is not hyperelliptic.
@@
Suppose by contradiction that $C$ is hyperelliptic. We mainly follow the reasoning in the introduction of [22]. By Lemma 7.1, $C$ is also bielliptic. Then, it is of the form $y^{2}=x^{8}+ax^{6}+bx^{4}+cx^{2}+1=f(x^{2})$, with $a,b,c\in\mathbb{F}_{q}$. Here $f$ is a polynomial of degree $4$. In this case, the double cover is given by the involution $(x,y)\mapsto(-x,y)$ and the elliptic curve is given by $y^{2}=f(x)$. Furthermore, quotienting $C$ by the hyperelliptic involution $(x,y)\mapsto(-x,-y)$ gives the genus $2$ curve $F\colon y^{2}=x\cdot f(x)$. Finally, we have $\mathop{\rm Jac}\nolimits(C)\sim E\times\mathop{\rm Jac}\nolimits(F)$, hence $\mathop{\rm Jac}\nolimits(F)\sim A$ by Lemma 7.1. The Jacobian of $F$ can be simple or split in the product of two elliptic curves. In both cases, this leads to a contradiction. ∎
@@
Let $q$ be the power of an odd prime. Let $C$ be an absolutely irreducible smooth genus $3$ curve lying on an abelian surface defined over $\mathbb{F}_{q}$ respecting one of the equivalent conditions of Theorem 1.2. Then, $C$ is a plane quartic of the form $y^{4}-h(x,z)y^{2}+r(x,z)=0$, where $h$ and $r$ are homogenous polynomials of degree $2$ and $4$, respectively, and one of the followings holds:
  1. (1)
@@
By Lemma 7.2 we know that $C$ cannot be hyperelliptic, and by hypothesis $C$ is the double cover of an elliptic curve $E$. Then, we can assume that the involution giving the double cover is $(x:y:z)\mapsto(x:-y:z)$, and therefore that $C$ can be written in the form $y^{4}-h(x,z)y^{2}+r(x,z)=0$, where $h$ is a homogenous degree $2$ polynomial, and $r$ is homogenous of degree $4$. Let us suppose that the polynomial $r$ has two quadratic factor $f$ and $g$ and that the polynomial $h(x,z)$ is not a linear combination of $f$ and $g$. Then, by [22, Thm. 1.1] we can explicitly construct a genus $2$ curve $F$ such that $\mathop{\rm Jac}\nolimits(C)\sim\mathop{\rm Jac}\nolimits(F)\times E$. Then, the curve $C$ cannot lie on $A$, since otherwise $A$ would be isogenous to $\mathop{\rm Jac}\nolimits(F)$ by Lemma 7.1. Finally, one of the two conditions in the proposition must hold. ∎
Finally, we present bounds on the number of rational points that a genus $3$ curve lying on an abelian surface with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$ can have. To start with, we recall and extend a result from [9] on the number of rational points on curves over abelian surfaces.
@@
Let $A$ be an abelian surface defined over $\mathbb{F}_{q}$ with trace $-a$. Let $C$ be an absolutely irreducible curve defined over $\mathbb{F}_{q}$, of arithmetic genus $p_{a}$, lying on $A$. Then
#$\mid\#C(\mathbb{F}_{q})-(q+1+a)\mid\leq|p_{a}-2|\lfloor 2\sqrt{q}\rfloor.$
@@
Note that in [9, Thm. 4] the author needs the hypothesis $a\geq-q$, which is however only used to prove the result for curves which are irreducible but not absolutely irreducible. For such curves, the number of rational points cannot exceed $p_{a}-1$, but we cannot prove a lower bound different from the trivial one.
@@
In particular, we see that an absolutely irreducible curve of genus $3$ lying on a simple abelian surface not isogenous to the Jacobian of a genus $2$ curve has not many rational points and it is far from reaching the Serre–Weil bound which states |#$|\#C(\mathbb{F}_{q})-(q+1)|\leq 3\lfloor 2\sqrt{q}\rfloor$.
@@ Acknowledgments
The first author would like to thank Christophe Ritzenthaler for preliminary discussions about the central question of this paper back in 2019, and for putting her in contact with the third author. The first and third authors thank Qing Liu for some useful discussion during the CAVARET conference. The authors are grateful to Jonas Bergström for comments on a preliminary version of the paper.
The first author is supported by the grant ANR-21-CE39-0009-BARRACUDA. The third author is supported by Nederlandse Organisatie voor Wetenschappelijk Onderzoek, grant number VI.Veni.202.107, and by Agence Nationale de la Recherche under the MELODIA project, grant number ANR-20-CE40-0013.

### v3

@@ 6. Computing isomorphism classes admitting a polarisation of degree $4$
In [19], it is described how to compute the isomorphism classes of abelian varieties over $\mathbb{F}_{q}$ belonging to an ordinary isogeny class $\mathcal{C}$ determined by an irreducible Weil polynomial $f(t)$. We summarise here the results that are relevant for us; see [19, Cor. 4.4, Thm. 5.4].
Let $K=\mathbb{Q}[t]/(f(t))=\mathbb{Q}[\pi]$ where $\pi$ denotes the class of $t$ in $K$ and set $R=\mathbb{Z}[\pi,q/\pi]$. There is an equivalence between the category of abelian varieties in $\mathcal{C}$ (with $\mathbb{F}_{q}$-morphisms) and the category of fractional $R$-ideals in $K$ (with $R$-linear morphisms). Hence, every overorder $S$ of $R$ occurs as the endomorphism ring of an abelian variety in $\mathcal{C}$. Moreover, the functor inducing the equivalence is compatible with duality and allows to describe polarisations as $R$-linear morphisms. For example, if $\mathop{\rm End}\nolimits(A)=S$ then $\mathop{\rm End}\nolimits(A^{\vee})=\overline{S}$. In [19, Sec. 6], the equivalence is used to produce an algorithm to compute the abelian varieties in $\mathcal{C}$ together with their polarisations (of a fixed degree) up to polarised isomorphism. We will use this algorithm in the next examples.
@@
Consider the isogeny class of abelian surfaces over $\mathbb{F}_{2}$ with label 2.2.a_ab determined by the Weil polynomial $f(t)=t^{4}-t^{2}+4$. According to the LMFDB, the class contains a Jacobian. Moreover, one computes that the order $\mathbb{Z}[\pi,4/\pi]$ has $3$ overorders $S_{1}$, $S_{2}$ and $\mathcal{O}$, where $\mathcal{O}$ is the maximal order of the number field $K=\mathbb{Q}[t]/(f(t))$. One observe that $S_{2}=\overline{S}_{1}$. Hence, each abelian surface $A$ with endomorphism ring $S_{1}$ cannot be isomorphic to its dual, which has endomorphism ring $S_{2}$. In particular, such an $A$ does not admit a principally polarisation. Hence, such an $A$ does not contain an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $2$ by Proposition 1.3 even if it is isogenous to a Jacobian; cf. Remark 1.5.
@@
In Theorem 5.7, we see that, in the isogeny class $\mathcal{C}$ associated to an ordinary Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, there is an abelian variety $A$ admitting a polarisation of degree $4$ if and only if there is an abelian variety $A^{\prime}$ in $\mathcal{C}$ with maximal endomorphism ring admitting a polarisation of degree $4$. This is not the case in general. For example, in the ordinary isogeny class with label 2.13.a_al determined by $t^{4}-11t^{2}+13^{2}$ and defined over $\mathbb{F}_{13}$ there are abelian varieties admitting a polarisation of degree $4$ but none of them has maximal endomorphism ring.
@@ 7. Genus $3$ curves lying on abelian surfaces
In this final section, we switch our attention from abelian surfaces to curves, and gather information on absolutely irreducible smooth genus $3$ curves lying on simple abelian surfaces. The following lemma characterises such curves as bielliptic.
@@
Let $q$ be the power of an odd prime. Let $C$ be an absolutely irreducible smooth genus $3$ curve lying on a simple abelian surface $A$, both defined over $\mathbb{F}_{q}$. Then there exists an elliptic curve $E$ defined over $\mathbb{F}_{q^{n}}$, with $n\leq 4$, such that $C$ is a double $\mathbb{F}_{q^{n}}$-cover of $E$.
@@
Here we mainly follow [5, Sec. 1.2], with some minor modifications since we are working over finite fields instead of over $\mathbb{C}$. By Theorem 2.2, the curve $C$ induces a polarisation $\lambda_{\mathcal{L}}:A\to A^{\vee}$ of degree $4$, where $\mathcal{L}=\mathcal{L}(C)$ is the ample line bundle defined by $C$. Let $\mathcal{B}$ be the base locus of $\mathcal{L}$, that is, =$\mathcal{B}=\{P\in A\,\mid\,s(P)=0\text{ for any section }s\in H^{0}(A,\mathcal{L})\}$ [10, II, Lem. 7.8]. Fix some $P_{0}\in\mathcal{B}$. Since $\ker\lambda_{\mathcal{L}}=\{P\in A\,\mid\,t_{P}^{*}\mathcal{L}\simeq\mathcal{L}\}$ [21, Ch. I, Sec. 8] we have, for any and $P\in\ker\lambda_{\mathcal{L}}$, $s(P_{0}+P)=t_{P}^{*}s(P_{0})=s(P_{0})=0$. Thus $P_{0}+\ker\lambda_{\mathcal{L}}\subseteq\mathcal{B}$. Since we are in characteristic different from $2$, $\ker\lambda_{\mathcal{L}}$ has order equal to $4$, the degree of $\lambda_{\mathcal{L}}$. On the other hand, $\mathcal{B}$ contains at most $C.C=4$ points. Hence, $P_{0}+\ker\lambda_{\mathcal{L}}$ and $\mathcal{B}$ coincide as sets of geometric points. Consider now the involution $\iota:A\to A$ given by $P\mapsto-P+2P_{0}$. Since the points in $\ker\lambda_{\mathcal{L}}$ are $2$-torsion points, the points in $\mathcal{B}$ are fixed points of $\iota$. Furthermore, we have $\iota^{*}\mathcal{L}\simeq\mathcal{L}$, and in particular $\iota C=C$ [5, Prop. 1.6]. Then one can lift $\iota$ to $C$ as an involution, and since $C$ is smooth, $\iota$ has exactly four fixed points on $C$, namely the points in $\mathcal{B}$ [5, 1.7]. Then, by the genus formula, the quotient $E=C/\iota$ is an elliptic curve, that is, $C$ is a double cover of $E$. Note that $\iota$ is not necessarily an $\mathbb{F}_{q}$-rational map: this will depend on whether $\mathcal{B}$ contains a rational point or not, that is, whether one can take $P_{0}$ to be rational. However, the fact that $\mathcal{B}$ is rational as a set implies that it is a (necessarily disjoint) union of orbits under the action of the Frobenius endomorphism. A point is of degree $n$ if the size of its orbit is $n$. As the size of $\mathcal{B}$ is $4$, the degree of its points is at most $4$. Hence the point $P_{0}$ defining the involution $\iota$ will have degree at most $4$. So $\iota$ and $E$ are defined over an extension of $\mathbb{F}_{q}$ of degree at most $4$. ∎
For the rest of the section, we fix an abelian surface $A$ defined over $\mathbb{F}_{q}$ with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, that is, which is simple and not isogenous to a Jacobian, or equivalently, not containing absolutely irreducible curves of geometric genus 00, $1$ or $2$ by Theorem 1.4. We suppose that there is an absolutely irreducible smooth genus $3$ curve $C$ defined over $\mathbb{F}_{q}$ that lies on $A$, and we want to study that curve.
We recall that absolutely irreducible smooth genus $3$ curves are either hyperelliptic curves or plane quartics. The next lemma shows that, under some conditions, hyperelliptic genus $3$ curves cannot lie on our abelian surfaces.
@@
Let $q$ be the power of an odd prime. Let $C$ be an absolutely irreducible smooth genus $3$ curve lying on an abelian surface $A$ defined over $\mathbb{F}_{q}$ which is simple and not isogenous to a Jacobian. Suppose further that the elliptic curve doubly covered by $C$ from Lemma 7.1 is defined over $\mathbb{F}_{q}$. Then $C$ is not hyperelliptic.
@@
Suppose by contradiction that $C$ is hyperelliptic. We mainly follow the reasoning in the introduction of [24]. By Lemma 7.1 and our hypothesis, $C$ is also bielliptic over $\mathbb{F}_{q}$. Then, it is of the form $y^{2}=x^{8}+ax^{6}+bx^{4}+cx^{2}+1=f(x^{2})$, with $a,b,c\in\mathbb{F}_{q}$. Here $f$ is a polynomial of degree $4$. In this case, the double cover is given by the involution $(x,y)\mapsto(-x,y)$ and the elliptic curve is given by $y^{2}=f(x)$. Furthermore, quotienting $C$ by the hyperelliptic involution $(x,y)\mapsto(-x,-y)$ gives the genus $2$ curve $F\colon y^{2}=x\cdot f(x)$. Finally, we have $\mathop{\rm Jac}\nolimits(C)\sim E\times\mathop{\rm Jac}\nolimits(F)$. On the other hand, as $C$ lies on $A$, we know by [9, Prop. 2]3 that $A$ is an isogeny factor of $\mathop{\rm Jac}\nolimits(C)$, that is, there exists an elliptic curve $E^{\prime}$ such that $\mathop{\rm Jac}\nolimits(C)\sim A\times E^{\prime}$. Since $A$ is simple we necessarily have $E\sim E^{\prime}$, and hence $\mathop{\rm Jac}\nolimits(F)\sim A$. The Jacobian of $F$ can be simple or split in the product of two elliptic curves. In both cases, this leads to a contradiction. ∎
@@
Let $q$ be the power of an odd prime. Let $C$ be an absolutely irreducible smooth genus $3$ curve lying on an abelian surface defined over $\mathbb{F}_{q}$ respecting one of the equivalent conditions of Theorem 1.4. Suppose further that the elliptic curve doubly covered by $C$ from Lemma 7.1 is defined over $\mathbb{F}_{q}$. Then, $C$ is a plane quartic of the form $y^{4}-h(x,z)y^{2}+r(x,z)=0$, where $h$ and $r$ are homogenous polynomials of degree $2$ and $4$, respectively, and one of the followings holds:
  1. (i)
@@
By Lemma 7.2 we know that $C$ cannot be hyperelliptic. Applying Lemma 7.1 and using the hypothesis of the proposition, $C$ is the double cover of an elliptic curve $E$ defined over $\mathbb{F}_{q}$. Then, we can assume that the involution giving the double cover is $(x:y:z)\mapsto(x:-y:z)$, and therefore that $C$ can be written in the form $y^{4}-h(x,z)y^{2}+r(x,z)=0$, where $h$ is a homogenous degree $2$ polynomial, and $r$ is homogenous of degree $4$. Let us suppose that the polynomial $r$ has two quadratic factor $f$ and $g$ and that the polynomial $h(x,z)$ is not a linear combination of $f$ and $g$. Then, by [24, Thm. 1.1] we can explicitly construct an absolutely irreducible smooth genus $2$ curve $F$ such that $\mathop{\rm Jac}\nolimits(C)\sim\mathop{\rm Jac}\nolimits(F)\times E$. Then, the curve $C$ cannot lie on $A$, since otherwise $A$ would be isogenous to $\mathop{\rm Jac}\nolimits(F)$ by Lemma 7.1. Finally, one of the two conditions in the proposition must hold. ∎
Finally, we present bounds on the number of rational points that an absolutely irreducible smooth genus $3$ curve lying on an abelian surface with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$ can have. To start with, we recall and extend a result from [9] on the number of rational points on curves over abelian surfaces.
@@
Let $A$ be an abelian surface defined over $\mathbb{F}_{q}$ with trace $-a$, that is, with Weil polynomial $t^{4}+at^{3}+bt^{2}+qat+q^{2}$. Let $C$ be an absolutely irreducible curve defined over $\mathbb{F}_{q}$, of arithmetic genus $p_{a}(C)=p_{a}$, lying on $A$. Then
|#$|\#C(\mathbb{F}_{q})-(q+1+a)|\leq|p_{a}-2|\lfloor 2\sqrt{q}\rfloor.$
@@
Note that in [9, Thm. 4] the author needs the hypothesis $a\geq-q$, which is however only used to prove the result for curves which are $\mathbb{F}_{q}$-irreducible but not absolutely irreducible. For such curves, the number of rational points cannot exceed $p_{a}-1$, but we cannot prove a lower bound different from the trivial one.
@@
In particular, we see that an absolutely irreducible curve of genus $3$ lying on a simple abelian surface not isogenous to the Jacobian of an absolutely irreducible smooth genus $2$ curve is far from reaching the Serre–Weil bound which states |#$|\#C(\mathbb{F}_{q})-(q+1)|\leq 3\lfloor 2\sqrt{q}\rfloor$.
@@ Acknowledgments
The first author would like to thank Christophe Ritzenthaler for preliminary discussions about the central question of this paper back in 2019, and for putting her in contact with the third author. The first and third authors thank Qing Liu for some useful discussion about curves on abelian surfaces during the CAVARET conference. The third author thanks Elisa Lorenzo García for valuable conversations. The authors are grateful to Jonas Bergström and Gaetan Bisson for comments on a preliminary version of the paper. Finally, the authors appreciate the valuable insights provided by Jean Kieffer and Damien Robert during the CAIPI symposium in Nancy, which led to the final version of Section 7.
The first author is supported by the grants ANR-21-CE39-0009-BARRACUDA and ANR-22-CPJ2-0047-01. The third author is supported by Nederlandse Organisatie voor Wetenschappelijk Onderzoek, grant number VI.Veni.202.107, by Agence Nationale de la Recherche under the MELODIA project, grant number ANR-20-CE40-0013, and by Marie-Skłodowska-Curie Actions - Postdoctoral Fellowships 2023 (project 101149209 - AbVarFq).